### v1

@@ V Conclusion
In conclusion, we proposed an algebraic construction defining a family of CSS quantum error-correcting codes, whose main advantage is a highly redundant set of minimum-weight stabilizer generators in codes with relatively short block lengths. Just like the higher-dimensional QHP codes which our codes generalize, the construction defines a sequence of chain complexes of varying lengths. The constructed $D$-dimensional codes are highly-symmetric and they are characterized by confinement profiles and single-shot properties similar to those of the $D$-dimensional QHP codes.
It is instructive to compare the parameters of $D$-dimensional AMC codes over $F[C_{\ell}]$ with a characteristic polynomial $h(x)$, and $D$-dimensional QHP codes constructed from a linear cyclic code with parameters $[\ell,\kappa,\delta]$ and the same check polynomial $h(x)$. At level $j$, the QHP code has block length $N={D\choose j}\ell^{D}$, dimension $K={D\choose j}\kappa^{D}$, and minimal distance $D=\min(\delta^{j},\delta^{D-j})$, i.e., its rate scales as power of that of the original code, $R=(\kappa/\ell)^{D}$. In comparison, the block lengths and the dimensions of AMC codes scale linearly with $\ell$, $n={D\choose j}\ell$ and $k={D\choose j}\kappa$, so that the rate does not change, which is also a feature of GB codes.
@@
Another nice property of the proposed construction is the high redundancy of minimum-weight stabilizer generators. While for a two-block code of dimension $k=2\kappa$ there are exactly $k$ minimum-weight redundant generators of each kind, stabilizer generator matrices of a balanced AMC code with $D=2j$ have ranks (${2j-1\choose j-1}(\ell-\kappa)$ and (${2j\choose j-1}\ell$ rows, which gives over a quarter redundant rows for $D=4$ and over a third for $D=6$. In particular, this redundancy can be used to speed up the decoding and improve its accuracy in a realistic setting of circuit noise. This is illustrated by our circuit simulations with relatively short AMC codes, which show two-shot decoding almost as accurate as full-block decoding, and a (pseudo)threshold above $1\%$, higher than that for surface codes.
@@ Appendix A Proof of Statement 2
The elimination process goes the same way for larger-$D$ complexes. Due to symmetry, whenever a pair of matrices, say, $A$ and $D$, are located in positions $(x_{1},y_{1})$, $(x_{1},y_{2})$ (same column) and in $(x_{2},y_{3})$, $(x_{3},y_{3})$ (same row), some other matrix, e.g., $B$, the $3\times 3$ block selected (rows $\{y_{1},y_{2},y_{3}\}$, columns $\{x_{1},x_{2},x_{3}\}$) will necessarily have the form of $Q_{2}$, up to some signs and the block labels.
@@
$=$ $\left[\begin{array}[c]{cccc}U&&&-V\\ &I\\ &&I\\ D^{\prime}&&&A^{\prime}\end{array}\right]\left[\begin{array}[c]{ccc|ccc}C&B&A\\ \hline\cr-D&&&B&A&0\\ &-D&&-C&0&A\\ &&-D&0&-C&-B\\ \end{array}\right]\ldots$
@@
$\mathop{\rm rank}R_{j}^{D+1}$ $=$ (${D-1\choose j-2}(\ell-\kappa)+{D-1\choose j-1}(\ell-\kappa)$
$=$ (${D\choose j-1}(\ell-\kappa),$
@@ Appendix B Proof of Statement 3
Permute the terms in the product so that $A$ enters last, and use the equations (33), (36), and (39). The block $N=A$ enters $Q_{j}$ only along the diagonal of the square block of block-dimension (${D-1\choose j-1}$; the remaining matrices are above and to the right. Using the assumed factorization, $B=B^{\prime}A$, etc., row and column transformation can be used to diagonalize the matrix $Q_{j}$, leaving only the square block with $A$ along the diagonal. This gives $\mathop{\rm rank}Q_{j}={D-1\choose j-1}\mathop{\rm rank}A$. ∎
@@
Now, we know that $Q_{j}(\boldsymbol{0})$ is identically zero. Consider $Q_{j}(\boldsymbol{\alpha})$ for $\boldsymbol{\alpha}\neq 0$. Without limiting generality, $\alpha_{1}\neq 0$. Due to the existence of the pseudoinverse $\underline{A}$, for every other (e.g., originating from the matrix $B$) we have
$E_{\boldsymbol{\alpha}}B=E_{\boldsymbol{\alpha}}E_{A}A\underline{A}B=E_{\boldsymbol{\alpha}}A\,\underline{A}B.$
This expression is valid even if $E_{\boldsymbol{\alpha}}A=0$, in which case $E_{\boldsymbol{\alpha}}=0$. Evidently, the projected blocks satisfy the conditions of the Lemma 6, which gives
@@
$=$ (${D-1\choose j-1}[\mathop{\rm rank}I_{\ell}-\mathop{\rm rank}E_{\boldsymbol{0}}]$
$=$ (${D-1\choose j-1}(\ell-\kappa),\quad\kappa\equiv\mathop{\rm rank}E_{\boldsymbol{0}},$
@@ Appendix D Proof of Statement 5
It is trivial to verify that the Statement is true for $j=0$ and for $j=D$. Denote $r\equiv m_{j}^{D}={D\choose j}$ and $c\equiv m_{j+1}^{D}={D\choose j+1}$ be the block dimensions of subspaces ${\cal Q}_{j}$ and ${\cal Q}_{j+1}$, respectively. For any non-zero vector $\boldsymbol{u}$ in $C_{h}^{\perp}$, we can construct $r$ same-weight vectors $\boldsymbol{u}_{i}$ in the space ${\cal Q}_{j}$ (with non-zero elements in only one block, and that equal to $\boldsymbol{x}$) which satisfy the equation $Q_{j}\boldsymbol{u}_{i}=\boldsymbol{0}$. We would like to show that all of them cannot be homologically trivial at the same time.
To this end, define $g(x)=(x^{\ell}-1)/h(x)$ and write the non-zero codeword as $u(x)=i(x)g(x)$, with $\deg i(x)<k$, and assume all $r$ vectors $\boldsymbol{u}_{i}$ to be trivial. This can be written as the matrix-polynomial equation
@@
where $\Xi_{r\times c}(x)$ is a matrix of specified dimensions over $F[x]/(x^{\ell}-1)$ which solves the equation. With the help of the invertible transformations similar to those constructed in Sec. A, matrix $Q_{j}(x)$ can be transformed to a diagonal form,
@@
with square matrices $W_{L}(x)$ and $W_{R}(x)$ invertible in the ring $F[x]/(x^{\ell}-1)$, and $\Lambda_{r\times c}$ with only
@@
non-zero elements (equal to 1), all in different rows and columns. Using the fact that the object on the left-hand-side (l.h.s.) is proportional to the identity matrix, we have
@@
$u(x)I_{r\times r}$ $=$ $h(x)\Lambda_{r\times c}{\Gamma_{r\times c}(x)}\bmod x^{\ell}-1,$
where the coefficients in $\Xi(x)$ have been replaced with the coefficients in $\Gamma(x)\equiv W_{R}(x)\,\Xi_{r\times c}(x)\,W_{L}(x)$; this transformation is one-to-one since the transformation matrices are invertible by construction.
If we pick a zero column in $\Lambda_{r\times c}$, the equation reads $u(x)=0\bmod x^{\ell}-1$, contradicting the initial assumption that $u(x)\neq 0$. Thus, for any non-zero codeword in $C_{h}^{\perp}$, there is a non-trivial vector in $H_{j}(\mathcal{Q})$ of the same weight, which proves the stated upper bound on the distance.∎

### v2

@@ V Conclusion
In conclusion, we constructed a family of CSS quantum error-correcting codes whose main advantage is a highly redundant set of minimum-weight stabilizer generators, in codes with relatively short block lengths. Just like the higher-dimensional QHP codes which our codes generalize, the construction defines a sequence of chain complexes of varying lengths. The constructed $D$-dimensional codes are highly-symmetric and they are characterized by confinement profiles and single-shot properties similar to those of the $D$-dimensional QHP codes.
It is instructive to compare the parameters of $D$-dimensional AMC codes over $F[C_{\ell}]$ with a characteristic polynomial $h(x)$, and $D$-dimensional QHP codes constructed from a linear cyclic code with parameters $\ell,\kappa,\delta]$ and the same check polynomial $h(x)$. At level $j$, the QHP code has block length $N={D\choose j}\ell^{D}$, dimension $K={D\choose j}\kappa^{D}$, and minimal distance $D=\min(\delta^{j},\delta^{D-j})$, i.e., its rate scales as power of that of the original code, $R=(\kappa/\ell)^{D}$. In comparison, the block lengths and the dimensions of AMC codes scale linearly with $\ell$, $n={D\choose j}\ell$ and $k={D\choose j}\kappa$, so that the rate does not change, which is also a feature of GB codes.
@@
Another nice property of the proposed construction is the high redundancy of minimum-weight stabilizer generators. While for a quasi-abelian two-block code of dimension $k=2\kappa$ there are exactly $\kappa$ minimum-weight redundant generators ($\kappa$ of each $X$ and $Z$ type), stabilizer generator matrices of a symmetric AMC code with even $D=2j$ have ranks ${2j-1\choose j-1}(\ell-\kappa)$ and $2j\choose j-1$$\ell$ rows, which gives over a quarter redundant rows for $D=4$ and over a third for $D=6$. In particular, this redundancy can be used to speed up the decoding and improve its accuracy in a realistic setting of circuit noise. This is illustrated by our circuit simulations with relatively short AMC codes, which show two-shot decoding almost as accurate as full-block decoding, and a (pseudo)threshold above $1\%$, higher than that for surface codes with weight-$4$ generators, or GB codes with stabilizer generators of weight $6$, same as for our AMC codes.
@@ Appendix A Proof of Statement 2
The elimination process goes the same way for larger-$D$ complexes. Due to symmetry, whenever a pair of matrices, say, $A$ and $B$, are located in positions $(x_{1},y_{1})$, $(x_{1},y_{2})$ (same column) and in $(x_{2},y_{3})$, $(x_{3},y_{3})$ (same row), the corresponding $3\times 3$ block (rows $\{y_{1},y_{2},y_{3}\}$, columns $\{x_{1},x_{2},x_{3}\}$) will necessarily have the form of $Q_{2}$ in Eq. (31), with only one additional block (e.g., $C$), up to some signs and the block labels.
@@
$=$ $\left[\begin{array}[c]{cccc}U&&&-V\\ &I\\ &&I\\ D^{\prime}&&&A^{\prime}\end{array}\right]\left[\begin{array}[c]{ccc|ccc}C&B&A\\ \hline\cr-D&&&B&A&0\\ &-D&&-C&0&A\\ &&-D&0&-C&-B\\ \end{array}\right]\overline{W}^{-}_{5,1}\overline{W}^{-}_{6,2}$
@@
$\mathop{\rm rank}R_{j}^{D+1}$ $=$ ${D-1\choose j-2}(\ell-\kappa)+{D-1\choose j-1}(\ell-\kappa)$
$=$ ${D\choose j-1}(\ell-\kappa),$
@@ Appendix B Proof of Statement 3
Permute the terms in the product so that $A$ enters last, and use the equations (33), (36), and (39). The block $N=A$ enters $Q_{j}$ only along the diagonal of the square block of block-dimension $D-1\choose j-1$}; the remaining matrices are above and to the right. Using the assumed factorization, $B=B^{\prime}A$, etc., row and column transformation can be used to diagonalize the matrix $Q_{j}$, leaving only the square block with $A$ along the diagonal. This gives $\mathop{\rm rank}Q_{j}={D-1\choose j-1}\mathop{\rm rank}A$. ∎
@@
Now, consider $Q_{j}(\boldsymbol{\alpha})$ for $\boldsymbol{\alpha}\neq 0$; without limiting generality, assume $\alpha_{1}\neq 0$, so that the matrix block $E_{\boldsymbol{\alpha}}A\neq 0$ has the same rank as the corresponding idempotent $E_{\boldsymbol{\alpha}}$. Due to the existence of the pseudoinverse $\underline{A}$, for every other matrix block, e.g., originating from the matrix $B$, we have
$E_{\boldsymbol{\alpha}}B=E_{\boldsymbol{\alpha}}E_{A}A\underline{A}B=(E_{\boldsymbol{\alpha}}A)\,\underline{A}B.$
Evidently, thus projected blocks contain factors $E_{\boldsymbol{\alpha}}A$ and thus satisfy the conditions of the Lemma 6, which gives
@@
$=$ ${D-1\choose j-1}[\mathop{\rm rank}I_{\ell}-\mathop{\rm rank}E_{\boldsymbol{0}}]$
$=$ ${D-1\choose j-1}(\ell-\kappa),\quad\kappa\equiv\mathop{\rm rank}E_{\boldsymbol{0}},$
@@ Appendix D Proof of Statement 5
It is trivial to verify that the Statement is true for $j=0$ and for $j=D$. Denote $r\equiv m_{j}^{D}={D\choose j}$ and $c\equiv m_{j+1}^{D}={D\choose j+1}$ the block dimensions of subspaces ${\cal Q}_{j}$ and ${\cal Q}_{j+1}$, respectively. For any non-zero vector $\boldsymbol{u}$ in $C_{h}^{\perp}$, we can construct $r$ same-weight vectors $\boldsymbol{u}_{i}$ in the space ${\cal Q}_{j}$ (with non-zero elements in only one block, and that equal to $\boldsymbol{u}$) which satisfy the equation $Q_{j}\boldsymbol{u}_{i}=\boldsymbol{0}$. We would like to show that all of them cannot be homologically trivial at the same time.
To this end, define $g(x)=(x^{\ell}-1)/h(x)$ and write the non-zero codeword as $u(x)=i(x)g(x)$, with $\deg i(x)<k$, and assume all $r$ vectors $\boldsymbol{u}_{i}$ to be trivial. The corresponding condition can be written as a matrix-polynomial equation
@@
where $\Xi_{r\times c}(x)$ is a matrix of unknowns of specified dimensions, with elements in the ring $F[x]/(x^{\ell}-1)$. With the help of the invertible transformations similar to those constructed in Sec. A, matrix $Q_{j}(x)$ can be transformed to a diagonal form,
@@
with square matrices $W_{L}(x)$ and $W_{R}(x)$ invertible in $F[x]/(x^{\ell}-1)$, and $\Lambda_{r\times c}$ with only
@@
non-zero elements (equal to 1), all in different rows and columns. Using the fact that the object on the left-hand-side (l.h.s.) is proportional to the identity matrix, after a multiplication by $W_{L}^{-1}(x)$ on the left and $W_{L}(x)$ on the right, we have
@@
$u(x)I_{r\times r}$ $=$ $h(x)\Lambda_{r\times c}{\Xi^{\prime}_{r\times c}(x)}\bmod x^{\ell}-1,$
where the arbitrary unknown coefficients in $\Xi(x)$ have been replaced with the unknown coefficients in $\Xi^{\prime}(x)\equiv W_{R}(x)\,\Xi_{r\times c}(x)\,W_{L}(x)$; this transformation is one-to-one since the transformation matrices are invertible by construction.
If we pick a zero column in $\Lambda_{r\times c}$, the equation reads $u(x)=0\bmod x^{\ell}-1$, contradicting the initial assumption. Thus, for any non-zero codeword in $C_{h}^{\perp}$, there is a non-trivial vector in $H_{j}(\mathcal{Q})$ of the same weight, which proves the stated upper bound on the distance.∎